# MORPH: PDE Foundation Models with Arbitrary Data Modality

### v3

@@ A.4 Data Processing
We use Reversible Instance Normalization (ReVIN) to counter covariate shift across heterogeneous datasets and training regimes. It was first proposed for time-series forecasting and validated to mitigate distribution shift (kim2021reversible). We normalize the data and cache the corresponding means and standard deviations, then use these statistics to exactly denormalize model outputs. Unlike mccabe2024multiple, which normalizes on-the-fly, we pre-normalize the entire dataset and train on normalized batches, reducing the normalization overhead incurred during training and fine-tuning, and inference.
@@
Input: Ordered files FILES, autoregressive window ar_order, chunk size, world size $W$, rank, num workers $K$, worker_i base_seed, epoch Output: Stream of $(X,y)$ pairs for this sub-worker
Count samples.
Compute $T=\sum\limits_{f\in\texttt{FILES}}N_{\text{sims}}(f)\cdot\max(0,\,T_{\text{steps}}(f)-\texttt{ar\_order})$;
Balance across sub-workers.
$G\leftarrow W\cdot K$; $T^{\ast}\leftarrow T-(T\bmod G)$; $m\leftarrow T^{\ast}/G$;
my_id $\leftarrow\texttt{rank}\cdot K+\texttt{worker\_id}$; global_idx $\leftarrow 0$;
Seed RNG $\leftarrow$ base_seed ++ epoch;
Stream and assign.;
foreach $f\in\texttt{FILES}$ do
   Open $f$ and iterate in chunks of size chunk_size;
   For each chunk: shuffle with the epoch seed and prepare AR pairs $(X,y)$;
 foreach $(x_{i},y_{i})\in(X,y)$ do
    if global_idx $\geq T^{\ast}$ then
       stop
    if (global_idx $\bmod G$) == my_id then
       yield $(x_{i},y_{i})$; if this sub-worker has yielded $m$ samples then
          stop
    global_idx $\leftarrow\texttt{global\_idx}+1$;
@@ B.1 Model details
Positional encodings: Let a learned absolute table $\mathbf{P}\in\mathbb{R}^{\texttt{max\_ar}\times\texttt{max\_patches}\times E}$ provide time–patch embeddings, and let token tensors be $x\in\mathbb{R}^{B\times t\times n\times E}$. At run time we construct $\widehat{\mathbf{P}}\in\mathbb{R}^{t\times n\times E}$ to match the current autoregressive context $t$ and patch count $n$, and add it to $x$ (broadcast over the batch). We use two different types of positional encoding.
ST-Bilinear: We resample $\mathbf{P}$ jointly over time and patches using bilinear interpolation:
@@
with interpolation weights $w^{(t)}$ and $w^{(n)}$ derived from the continuous reindexing of $1,\texttt{max\_ar}]\to[1,t]$ and $1,\texttt{max\_patches}]\to[1,n]$ (no temporal slicing). This preserves smooth variation across both axes and supports arbitrary $(t,n)$ without changing parameter count. We employ this variant for the L model with $\texttt{max\_ar}=16$.
@@
yielding $\widehat{\mathbf{P}}\in\mathbb{R}^{t\times n\times E}$. This avoids temporal smoothing, preserves the semantics of discrete AR time steps, and is lighter-weight computationally. We use this variant for the Ti/S/M models with $\texttt{max\_ar}=1$.
Both encodings are absolute and enable variable $(t,n)$ by resampling, with an $\mathcal{O}(t\,n\,E)$ application cost. ST-Bilinear provides smooth extrapolation in time and space—beneficial for long AR horizons (L model). S-Linear/T-Slice enforces a strict AR horizon ($t\leq\texttt{max\_ar}$) and avoids temporal aliasing—well-suited to short-context regimes (Ti/S/M). We apply dropout after constructing $\widehat{\mathbf{P}}$ and then add it to token embeddings prior to attention.
@@ B.2 Pretraining settings
Low-Rank Adaptation (LoRA) is used for fine-tuning LLMs where the pretrained model weights are frozen while introducing trainable rank decomposition matrices into the transformer layers, dramatically reducing task-specific parameters during finetuning (hu2022lora). The motivation is supported by evidence that over-parameterized networks possess low intrinsic dimensionality, suggesting that task-specific updates lie in a low-rank subspace (aghajanyan2020intrinsic). Mathematically, for a frozen weight $W_{0}\in\mathbb{R}^{d\times k}$, LoRA parameterizes the update as
@@
  • Optimizer: We use AdamW (loshchilov2017decoupled) with a learning rate of $10^{-3}$ and weight decay $10^{-2}$. For standalone CFD datasets, we use $10^{-4}$. The learning-rate schedule is ReduceLROnPlateau with a decay factor of $0.5$ and a patience of $5$ epochs, applied after a $20$-epoch warmup.
@@ B.3.2 Existing SOTA Models
To contextualize MORPH against recent PDE foundation models, we fine-tune two Poseidon variants (Poseidon-T (21M) and Poseidon-B (158M) (herde2024poseidon)) and two DPOT variants (DPOT-S (30M) and DPOT-M (122M) (hao2024dpot)). Results are presented in Tables 5, 8 and 9.
@@
For the rollout experiment (Tables 8 and 9), we fine-tune on the first 128 trajectories for 200 epochs and evaluate on the last 240 trajectories of the FNS-KF subset of PDEBench, reporting average MSE and RMSE over 20-step rollouts (takamoto2022pdebench). Throughout, all inputs are at $128{\times}128$ spatial resolution per channel, and we apply the data normalization procedures specified by each repository.
@@
DPOT is pretrained with a 10-step temporal context at $128{\times}128$ resolution. To adapt DPOT for single-step $t\!\rightarrow\!t{+}1$ prediction (Table 5), we follow the input-padding strategy analogous to herde2024poseidon, feeding 10 copies of the state at time $t$ to satisfy the context window:
@@ C.2 Ablation studies
We pretrained MORPH-FM-S (30M) from scratch on the same pretraining corpus as Poseidon i.e., NS-Sines, NS-Gaussians, CE-RP, CE-CRP, CE-KH, CE-Gauss herde2024poseidon for 100 epochs. We employ a ReduceLROnPlateau scheduler with a wait of 5 epochs and early stopping with a patience of 10 epochs. We then finetune MORPH on 128 trajectories and test on 240 trajectories for 200 epochs across three downstream tasks: CE-RM, NS-PwC, and NS-SL. We use a learning rate of 1e-4, batch size 40, and weight decay 1e-4, early stopping with patience of 25 during the finetuning. The hyperparameter settings for Poseidon are presented in Appendix B.3.2. Table 8 shows the comparison. Full autoregressive rollouts are shown in Figs. 10, 12 and 11.
Table 9: Ablation studies (Pretraining corpus): MORPH-FM-S*(27M) and Poseidon-FM-B are finetuned with 128 trajectories for 200 epochs: Rollout MSE on test set (lower is better) average over 240 trajectories.
@@ C.3.1 Standalone surrogate
Figs. 6 and 7 show 10-step autoregressive rollouts for MORPH-SS-Ti and MORPH-SS-S on the Shallow Water Equations (SWE) dataset, using the t=0t{=}0 snapshot as input. Across the rollout horizon, MORPH-SS-S exhibits lower error than MORPH-SS-Ti, consistent with the NRMSE reported in Table 3. Both models produce stable multi-step forecasts, exhibiting limited error accumulation and no blow-up for the 10-step horizon.
@@ C.3.2 Finetuned foundation models
Figs. 8 and 9 present 10-step autoregressive (AR) rollouts for MORPH-FM-Ti and MORPH-FM-S fine-tuned on the forced incompressible Navier–Stokes with Kolmogorov forcing (FNS-KF) dataset, initialized from the t=0 snapshot. Over the 10-step horizon, MORPH-FM-S achieves consistently lower error than MORPH-FM-Ti. Both models remain stable under rollouts with limited error accumulation and no concerning instability across the rollout.
@@
Figure 9: Finetuning MORPH-FM-S ($\sim$ 30M) for FNS-KF prediction task: 10-step autoregressive rollouts for $v_{x}$ with $t=0$ (initial frame) as input.
@@ C.3.3 Finetuned foundation model (Pretrained with PDEGym)
Figure 10: Full autoregressive rollouts for MORPH-FM-S*(27M) finetuned for NS-SL prediction task with 128 trajectories, tested on 240 trajectories for 200 epochs
@@
Figure 11: Full autoregressive rollouts for MORPH-FM-S*(27M) finetuned for NS-PwC prediction task with 128 trajectories, tested on 240 trajectories for 200 epochs
@@
Figure 12: Full autoregressive rollouts for MORPH-FM-S*(27M) finetuned for CE-RM prediction task with 128 trajectories, tested on 240 trajectories for 200 epochs

### v4

@@ A.5 Data Processing
We use Reversible Instance Normalization (ReVIN) to counter covariate shift across heterogeneous datasets and training regimes. It was first proposed for time-series forecasting and validated to mitigate distribution shift (Kim et al., 2021). We normalize the data and cache the corresponding means and standard deviations, then use these statistics to exactly denormalize model outputs. Unlike McCabe et al. (2024), which normalizes on-the-fly, we pre-normalize the entire dataset and train on normalized batches, reducing the normalization overhead incurred during training and fine-tuning, and inference.
@@
0: Ordered files FILES, autoregressive window ar_order, chunk size chunk_size, world size $W$, rank, num workers $K$, worker_id, base_seed, epoch
0: Stream of $(X,y)$ pairs for this sub-worker
[…]
7: Seed RNG $\leftarrow\texttt{base\_seed}+\texttt{epoch}$
8:for all $f\in\texttt{FILES}$ do
9:  Open $f$ and iterate in chunks of size chunk_size
10:  Shuffle each chunk with the epoch seed and prepare AR pairs $(X,y)$
11:  for all $(x_{i},y_{i})\in(X,y)$ do
12:   if $\texttt{global\_idx}\geq T^{\ast}$ then
13:    stop
14:   end if
15:   if $(\texttt{global\_idx}\bmod G)=\texttt{my\_id}$ then
16:    yield $(x_{i},y_{i})$
17:    if this sub-worker has yielded $m$ samples then
18:     stop
19:    end if
20:   end if
21:   $\texttt{global\_idx}\leftarrow\texttt{global\_idx}+1$
22:  end for
23:end for
@@ B.1 Model details
Positional encodings: Let a learned absolute table $\mathbf{P}\in\mathbb{R}^{\texttt{max\_ar}\times\texttt{max\_patches}\times E}$ provide time–patch embeddings, and let token tensors be $x\in\mathbb{R}^{B\times t\times n\times E}$. At runtime we construct $\widehat{\mathbf{P}}\in\mathbb{R}^{t\times n\times E}$ to match the current autoregressive context $t$ and patch count $n$, and add it to $x$ (broadcast over the batch). Here, “patches” refer to the flattened patch sequence index obtained from a fixed rasterization of the $(D,H,W)$ patch grid. We use two different types of learnable positional encoding.
ST-Bilinear: We resample $\mathbf{P}$ jointly over time and patches using separable bilinear interpolation (linear in each axis):
@@
with interpolation weights $w^{(t)}$ and $w^{(n)}$ given by the standard 1D linear interpolation induced by continuous reindexing of $1,\texttt{max\_ar}]\to[1,t]$ and $1,\texttt{max\_patches}]\to[1,n]$ (no temporal slicing; indices outside the source range are clamped). This preserves smooth variation across both axes and supports arbitrary $(t,n)$ without changing parameter count. We employ this variant for the L model with $\texttt{max\_ar}=16$.
@@
yielding $\widehat{\mathbf{P}}\in\mathbb{R}^{t\times n\times E}$. This avoids temporal smoothing, preserves the semantics of discrete AR time steps, and is lighter-weight computationally. We use this variant for the Ti/S/M models with $\texttt{max\_ar}=1$ (consistent with the $AR(1)$ setting used in this paper).
Both encodings are learnable and enable variable $(t,n)$ by resampling, with an $\mathcal{O}(t\,n\,E)$ application cost. ST-Bilinear provides smooth extrapolation in time and space which is beneficial for long AR horizons (L model). S-Linear/T-Slice enforces a strict AR horizon ($t\leq\texttt{max\_ar}$) and avoids temporal aliasing well-suited to short-context regimes (Ti/S/M). We apply dropout after constructing $\widehat{\mathbf{P}}$ and then add it to token embeddings prior to attention.
@@ B.2 Pretraining settings
Low-Rank Adaptation (LoRA) is used for fine-tuning LLMs where the pretrained model weights are frozen while introducing trainable rank decomposition matrices into the transformer layers, dramatically reducing task-specific parameters during finetuning (Hu et al., 2022). The motivation is supported by evidence that over-parameterized networks possess low intrinsic dimensionality, suggesting that task-specific updates lie in a low-rank subspace (Aghajanyan et al., 2020). Mathematically, for a frozen weight $W_{0}\in\mathbb{R}^{d\times k}$, LoRA parameterizes the update as
@@
  • Optimizer: We use AdamW (Loshchilov and Hutter, 2017) with a learning rate of $10^{-3}$ and weight decay $10^{-2}$. For standalone CFD datasets, we use $10^{-4}$. The learning-rate schedule is ReduceLROnPlateau with a decay factor of $0.5$ and a patience of $5$ epochs, applied after a $20$-epoch warmup.
@@ B.3.2 Existing SOTA Models
To contextualize MORPH against recent PDE foundation models, we fine-tune two Poseidon variants (Poseidon-T (21M) and Poseidon-B (158M) (Herde et al., 2024)) and two DPOT variants (DPOT-S (30M) and DPOT-M (122M) (Hao et al., 2024)). Results are presented in Tables 3, 8 and 9.
@@
For the rollout experiment (Tables 8 and 9), we fine-tune on the first 128 trajectories for 200 epochs and evaluate on the last 240 trajectories of the FNS-KF subset of PDEBench, reporting average MSE and RMSE over 20-step rollouts (Takamoto et al., 2022). Throughout, all inputs are at $128{\times}128$ spatial resolution per channel, and we apply the data normalization procedures specified by each repository.
@@
DPOT is pretrained with a 10-step temporal context at $128{\times}128$ resolution. To adapt DPOT for single-step $t\!\rightarrow\!t{+}1$ prediction (Table 3), we follow the input-padding strategy analogous to (Herde et al., 2024), feeding 10 copies of the state at time $t$ to satisfy the context window:
@@ C.2 Ablation studies
We pretrained MORPH-FM-S (30M) from scratch on the same pretraining corpus as Poseidon i.e., NS-Sines, NS-Gaussians, CE-RP, CE-CRP, CE-KH, CE-Gauss (Herde et al., 2024) for 100 epochs. We employ a ReduceLROnPlateau scheduler with a wait of 5 epochs and early stopping with a patience of 10 epochs. We then finetune MORPH on 128 trajectories and test on 240 trajectories for 200 epochs across three downstream tasks: CE-RM, NS-PwC, and NS-SL. We use a learning rate of 1e-4, batch size 40, and weight decay 1e-4, early stopping with patience of 25 during the finetuning. The hyperparameter settings for Poseidon are presented in Appendix B.3.2. Table 8 shows the comparison. Full autoregressive rollouts are shown in Figs. 9, 11 and 10.
Table 9: Ablation studies (Pretraining corpus): MORPH-FM-S*(27M) and Poseidon-FM-B are fine-tuned with 128 trajectories for 200 epochs: Rollout MSE on test set (lower is better) average over 240 trajectories.
@@ C.3.1 Standalone surrogate
Figs. 5 and 6 show 10-step autoregressive rollouts for MORPH-SS-Ti and MORPH-SS-S on the Shallow Water Equations (SWE) dataset, using the t=0t{=}0 snapshot as input. Across the rollout horizon, MORPH-SS-S exhibits lower error than MORPH-SS-Ti, consistent with the NRMSE reported in Table 1. Both models produce stable multi-step forecasts, exhibiting limited error accumulation and no blow-up for the 10-step horizon.
@@ C.3.2 Finetuned foundation models
Figs. 7 and 8 present 10-step autoregressive (AR) rollouts for MORPH-FM-Ti and MORPH-FM-S fine-tuned on the forced incompressible Navier–Stokes with Kolmogorov forcing (FNS-KF) dataset, initialized from the t=0 snapshot. Over the 10-step horizon, MORPH-FM-S achieves consistently lower error than MORPH-FM-Ti. Both models remain stable under rollouts with limited error accumulation and no concerning instability across the rollout.
@@
Figure 8: Finetuning MORPH-FM-S ($\sim$ 30M) for FNS-KF prediction task: 10-step autoregressive rollouts for $v_{x}$ and $v_{y}$ with $t=0$ (initial frame) as input.
@@ C.3.3 Finetuned foundation model (Pretrained with PDEGym)
Figure 9: Full autoregressive rollouts for MORPH-FM-S*(27M) fine-tuned for NS-SL prediction task with 128 trajectories, tested on 240 trajectories for 200 epochs
@@
Figure 10: Full autoregressive rollouts for MORPH-FM-S*(27M) fine-tuned for NS-PwC prediction task with 128 trajectories, tested on 240 trajectories for 200 epochs
@@
Figure 11: Full autoregressive rollouts for MORPH-FM-S*(27M) fine-tuned for CE-RM prediction task with 128 trajectories, tested on 240 trajectories for 200 epochs